# Cross-Modality Perturbation Synergy Attack for Person Re-identification

### v3

@@ Abstract
footnotetext: Min Jiang and Yunpeng Gong are with the Department of Artificial Intelligence, Key Laboratory of Multimedia Trusted Perception and Efficient Computing, Ministry of Education of China School of Informatics, Key Laboratory of Digital Protection and Intelligent Processing of Fujian and Taiwan Intangible Cultural Heritage, Ministry of Culture and Tourism, Xiamen University, Xiamen 361005, Fujian, P.R. China (e-mail: minjiang@xmu.edu.cn; gongyunpeng@stu.xmu.edu.cn or fmonkey625@gmail.com).
@@ 1 Introduction
Currently, most research on the security of ReID focuses on single-modality systems based on RGB images [47, 39, 6, 1, 2, 10, 29], while the security of cross-modality ReID systems has received insufficient attention. The challenge in cross-modality attacks arises from significant visual differences among different modality inputs, requiring attackers to effectively capture shared features from each modality for perturbation implementation. However, as shown in Fig. 1, existing attack methods in cross-modal scenarios require optimizing perturbations separately for each modality, lacking an intrinsic mechanism to capture shared knowledge between different modalities, which limits the success rate of the attacks. To address this issue, we propose a synergistic optimization method combined with triplet loss, utilizing information from different modalities to optimize the universal perturbation. This method pushes the features of different samples into a common sub-region that affects the model’s accuracy, as shown in Fig. 2.
@@
In our future work, we will consider incorporating evolutionary search algorithms [32, 14] to capture shared knowledge across different modalities, thereby enhancing the transferability of attacks in any given modality.
@@ 2 Related Works
Adversarial Attacks in ReID. Some ReID attack methods have been proposed, with current research predominantly focusing on RGB-RGB matching. These methods mainly include: Metric-FGSM [1] extends some techniques, inspired by classification attacks, into a category known as metric attacks. These encompass Fast Gradient Sign Method (FGSM) [11], Iterative FGSM (IFGSM), and Momentum IFGSM (MIFGSM) [4]. The Furthest-Negative Attack (FNA) [2] integrates hard sample mining [13] and triple loss to employ pushing and pulling guides. These guides guide image features towards the least similar cluster while moving away from other similar features. Deep Mis-Ranking (DMR) [29] utilizes a multi-stage network architecture to pyramidally extract features at different levels, aiming to derive general and transferable features for adversarial perturbations. Gong et al. [6] proposed a method specifically for attacking color features without requiring additional reference images and discussed effective defense strategies against current ReID attacks. The Opposite-direction Feature Attack (ODFA) [47] exploits feature-level adversarial gradients to generate examples that guide features in the opposite direction with an artificial guide. Yang et al. [39] introduced a combined attack named Col.+Del. (Color Attack and Delta Attack), which integrates UAP-Retrieval [16] with color space perturbations [15]. While this method also explores universal perturbations in ReID, its generality is limited due to the inability to leverage color information in cross-modality problems and the lack of a mechanism for associating different modality information. In contrast to the aforementioned approaches, our focus lies on addressing cross-modality challenges.
@@ 3.2 Optimizing Loss Functions for Attacking
As shown in Fig. 4, the loss function mentioned above fully leverages the triplet-wise relationships across different modality. Through this loss, we are able to pull the negative samples of each modality closer to the adversarial samples and push the positive samples of each modality away from the adversarial samples. Here, $C^{p}_{v}$ and $C^{n}_{v}$ represent the cluster centroids of the positive samples to push and negative samples to pull, respectively, in the original visible (RGB) image feature space of the training data. Similar definitions apply to other modalities. $f^{adv}_{v}$, $f^{adv}_{g}$, and $f^{adv}_{ir}$ denote the perturbed features of the disturbed image in the visible, grayscale, and infrared (or thermal) modalities, respectively.
@@ 3.4 Cross-Modality Perturbation Synergy Attack
$\quad\left.-\left.\lVert C^{p}_{g}-f^{adv}_{ir}\rVert_{2}+\rho\right),0\right]$
$\Delta_{ir}=\theta\Delta_{RGB}+\frac{\nabla_{\eta}L_{ir}}{\|\nabla_{\eta}L_{ir}\|_{1}}$ (8)
$\eta=\text{clip}(\eta+\alpha\cdot\text{sign}(\Delta_{IR}),-\varepsilon,\varepsilon)$ (9)
@@
Theoretical Analysis. In traditional optimization, after optimizing for one modality, the perturbation may become suboptimal for the other, resulting in higher errors and bias towards a single modality. In contrast, the proposed aggregated optimization method jointly optimizes both modalities by simultaneously considering their losses, ultimately finding a universal perturbation that improves cross-modality performance. In the supplementary material 7, we provide a mathematical analysis demonstrating the effectiveness of this method compared to traditional attack methods that lack intrinsic correlations between different modalities.
@@ 4 Experiments
In this section, we compare our approach with several methods, including traditional classification attack methods FGSM [11] and PGD [19], traditional metric attack methods like Metric-FGSM [1], as well as state-of-the-art ReID attack methods such as LTA [6] The LTA code is available at: https://github.com/finger-monkey/LTA_and_joint-defence , ODFA[47] and Col.+Del.[39]. Our experiments were conducted using three RTX 2080 Ti GPUs, each with 11GB of memory.
Datasets. We evaluate our proposed method on two commonly used cross-modality ReID datasets: SYSU-MM01 [34], RegDB [22] and LLCM [43]. SYSU-MM01 is a large-scale dataset with 395 training identities, captured by 6 cameras (4 RGB, 2 near-infrared) on the SYSU campus. It comprises 22,258 visible and 11,909 near-infrared images. The testing set consists of 95 identities with two evaluation settings. The query sets include 3803 images from two IR cameras. We conduct ten trials following established methods [33] and report the average retrieval performance. Please refer to [34] for the evaluation protocol. RegDB [22] is a smaller-scale dataset with 412 identities, each having ten visible and ten thermal images. we randomly select 206 identities (2,060 images) for training and use the remaining 206 identities (2,060 images) for testing. We assess our model in two retrieval scenarios: visible-thermal and thermal-visible performance. LLCM is a dataset designed specifically for cross-modality ReID in low-light environments. Compared to other datasets, its diverse scenarios and low-light conditions present greater challenges for attackers. This complexity and uncertainty make adversarial attacks more difficult to execute.
@@ 4.1 Performance on Cross-Modality ReID
From Tab.1, it can be seen that the proposed method reduces the rank-1 accuracy to below 2% in both the ’Visible to Infrared’ and ’Infrared to Visible’ cases. Similarly, from Tab.2, the rank-1 accuracy drops below 3% in both the ’Visible to Thermal’ and ’Thermal to Visible’ scenarios. In contrast, traditional metric-based attacks like Metric-FGSM (M-FGSM)[1] and ODFA[47] result in attacked models with rank-1 accuracy above 14%, while traditional classification attacks (such as FGSM [11] and PGD [19]) perform even worse, with rank-1 accuracy remaining over 60%. This is because ReID relies on metric learning for feature matching rather than category classification, requiring attacks specifically tailored for metric learning. These results indicate: 1) Compared to traditional methods that optimize perturbations separately for each modality without considering the inherent correlations between different modalities, our proposed approach demonstrates significant attacking effectiveness across different modalities. 2) From the perspective of attack effectiveness, cross-modality retrieval scenarios such as ’Thermal to Visible’ appear to be relatively less susceptible to attacks (with higher mean Average Precision than ’Infrared to Visible’). We speculate that this may be due to thermal infrared images having fewer texture details, reducing discernible features for attackers and thus lowering the success rate of attacks.
@@ 5 Conclusion
In this study, we have proposed a cross-modality attack method known as Cross-Modality Perturbation Synergy (CMPS) attack, aimed at evaluating the security of cross-modality ReID systems. The core idea behind the CMPS attack is to capture shared knowledge between visible and non-visible images to optimize perturbations. Additionally, we proposed a cross-modality attack augmentation method, utilizing grayscale images to bridge the gap between different modalities, further enhancing the attack performance. Through experiments conducted on the RegDB and SYSU datasets, we demonstrated the effectiveness of the proposed method while also revealing the limitations of traditional attack approaches. The primary objective of this study has been to assess the security of cross-modality ReID systems. In future research, we plan to develop robust ReID methods specifically tailored for cross-modality attacks, aimed at defending against adversarial samples. This study not only contributes to advancing the understanding of the security of cross-modality ReID systems but also provides strong motivation for ensuring the reliability and security of these systems in real-world applications.

### v4

@@ Abstract
**footnotetext: Min Jiang and Yunpeng Gong are with the Department of Artificial Intelligence, Key Laboratory of Multimedia Trusted Perception and Efficient Computing, Ministry of Education of China, School of Informatics, Key Laboratory of Digital Protection and Intelligent Processing of Intangible CulturalHeritage of Fujian and Taiwan, Ministry of Culture and Tourism, Xiamen University, Xiamen 361005, Fujian, P.R. China (e-mail: minjiang@xmu.edu.cn; gongyunpeng@stu.xmu.edu.cn or fmonkey625@gmail.com).
@@ 1 Introduction
Currently, most research on the security of ReID focuses on single-modality systems based on RGB images [23, 24, 25, 26, 27, 28], while the security of cross-modality ReID systems has received insufficient attention. The challenge in cross-modality attacks arises from significant visual differences among different modality inputs, requiring attackers to effectively capture shared features from each modality for perturbation implementation. However, as shown in Fig. 1, existing attack methods in cross-modal scenarios require optimizing perturbations separately for each modality, lacking an intrinsic mechanism to capture shared knowledge between different modalities, which limits the success rate of the attacks. To address this issue, we propose a synergistic optimization method combined with triplet loss, utilizing information from different modalities to optimize the universal perturbation. This method pushes the features of different samples into a common sub-region that affects the model’s accuracy, as shown in Fig. 2.
@@ 2 Related Works
Adversarial Attacks in ReID. Some ReID attack methods have been proposed, with current research predominantly focusing on RGB-RGB matching. These methods mainly include: Metric-FGSM [26] extends some techniques, inspired by classification attacks, into a category known as metric attacks. These encompass Fast Gradient Sign Method (FGSM) [13], Iterative FGSM (IFGSM), and Momentum IFGSM (MIFGSM) [39]. The Furthest-Negative Attack (FNA) [27] integrates hard sample mining [40] and triple loss to employ pushing and pulling guides. These guides guide image features towards the least similar cluster while moving away from other similar features. Deep Mis-Ranking (DMR) [28] utilizes a multi-stage network architecture to pyramidally extract features at different levels, aiming to derive general and transferable features for adversarial perturbations. Gong et al. [25] proposed a local transformation attack (LTA) method specifically aimed at attacking color features without requiring additional reference images, and discussed effective defense strategies against current ReID attacks. The Opposite-direction Feature Attack (ODFA) [23] exploits feature-level adversarial gradients to generate examples that guide features in the opposite direction with an artificial guide. Yang et al. [24] introduced a combined attack named Col.+Del., which integrates UAP-Retrieval [41] with color space perturbations [42]. While this method also explores universal perturbations in ReID, its generality is limited due to the inability to leverage color information in cross-modality problems and the lack of a mechanism for associating different modality information. In contrast to the aforementioned approaches, our focus lies on addressing cross-modality challenges.
@@ 3.2 Optimizing Loss Functions for Attacking
As shown in Fig. 4, the loss function mentioned above fully leverages the triplet-wise relationships across different modality. Through this loss, we are able to pull the negative samples of each modality closer to the adversarial samples and push the positive samples of each modality away from the adversarial samples. Here, $C^{p}_{RGB}$ and $C^{n}_{RGB}$ represent the cluster centroids of the positive samples to push and negative samples to pull, respectively, in the original visible (RGB) image feature space of the training data. Similar definitions apply to other modalities. $f^{adv}_{RGB}$, $f^{adv}_{g}$, and $f^{adv}_{ir}$ denote the perturbed features of the disturbed image in the visible, grayscale, and infrared (or thermal) modalities, respectively.
@@ 3.4 Cross-Modality Perturbation Synergy Attack
$\quad\left.-\left.\lVert C^{p}_{g}-f^{adv}_{ir}\rVert_{2}+\rho\right),0\right],$
$\Delta_{ir}=\theta\Delta_{RGB}+\frac{\nabla_{\eta}L_{ir}}{\|\nabla_{\eta}L_{ir}\|_{1}},$ (8)
$\eta=\text{clip}(\eta+\alpha\cdot\text{sign}(\Delta_{ir}),-\varepsilon,\varepsilon).$ (9)
@@
Theoretical Analysis. In traditional optimization, optimizing for one modality can render the perturbation suboptimal for the other, leading to a bias toward a single modality. In contrast, the proposed aggregated optimization method jointly optimizes both modalities, ultimately identifying a universal perturbation that enhances cross-modality attack performance. In the supplementary material 7, we provide a mathematical analysis demonstrating the effectiveness of this method compared to traditional attack methods that lack intrinsic correlations between different modalities.
@@ 4 Experiments
In this section, we compare our approach with several methods, including traditional classification attack methods FGSM [13] and PGD [14], traditional metric attack methods like Metric-FGSM [26], as well as state-of-the-art ReID attack methods such as LTA [25] ***The LTA code is available at: https://github.com/finger-monkey/LTA_and_joint-defence , ODFA[23] and Col.+Del.[24].
Datasets. We evaluate our proposed method on two commonly used cross-modality ReID datasets: SYSU-MM01 [30], RegDB [29] and LLCM [21]. SYSU-MM01 is a large-scale dataset with 395 training identities, captured by 6 cameras (4 RGB, 2 near-infrared) on the SYSU campus. It comprises 22,258 visible and 11,909 near-infrared images. The testing set consists of 95 identities with two evaluation settings. The query sets include 3803 images from two IR cameras. We conduct ten trials following established methods [43] and report the average retrieval performance. Please refer to [30] for the evaluation protocol. RegDB [29] is a smaller-scale dataset with 412 identities, each having ten visible and ten thermal images. we randomly select 206 identities (2,060 images) for training and use the remaining 206 identities (2,060 images) for testing. LLCM is a dataset designed specifically for cross-modality ReID in low-light environments. Compared to other datasets, its diverse scenarios and low-light conditions present greater challenges for attackers. This complexity and uncertainty make adversarial attacks more difficult to execute. We assess our model in two retrieval scenarios: visible-thermal and thermal-visible performance.
@@ 4.1 Performance on Cross-Modality ReID
From Tab.1, it can be seen that the proposed method reduces the rank-1 accuracy to below 2% in both the ’Visible to Infrared’ and ’Infrared to Visible’ cases. Similarly, from Tab.2, the rank-1 accuracy drops below 3% in both the ’Visible to Thermal’ and ’Thermal to Visible’ scenarios. In contrast, traditional metric-based attacks, such as Metric-FGSM (M-FGSM)[26], LTA [25] and ODFA[23], lead to attacked models with significantly higher rank-1 accuracy, whereas traditional classification attacks (such as FGSM [13] and PGD [14]) perform even worse, with rank-1 accuracy remaining over 60%. This is because ReID relies on metric learning for feature matching rather than category classification, requiring attacks specifically tailored for metric learning. These results indicate that, compared to traditional methods that optimize perturbations separately for each modality without considering the inherent correlations between different modalities, our proposed approach demonstrates significant attacking effectiveness across different modalities.
@@ 5 Conclusion
In this study, we have proposed a cross-modality attack method known as Cross-Modality Perturbation Synergy (CMPS) attack, aimed at evaluating the security of cross-modality ReID systems. The core idea behind the CMPS attack is to capture shared knowledge between visible and non-visible images to optimize perturbations. Additionally, we proposed a cross-modality attack augmentation method, utilizing grayscale images to bridge the gap between different modalities, further enhancing the attack performance. Through experiments conducted on the RegDB, SYSU and LLCM datasets, we demonstrated the effectiveness of the proposed method while also revealing the limitations of traditional attack approaches. The primary objective of this study has been to assess the security of cross-modality ReID systems. In future research, on the one hand, we will continue to improve the transferability of cross-modality attacks across different datasets and models; on the other hand, we plan to develop robust ReID methods specifically tailored for cross-modality attacks, aimed at defending against adversarial samples. This study not only contributes to advancing the understanding of the security of cross-modality ReID systems but also provides strong motivation for ensuring the reliability and security of these systems in real-world applications.
@@ 6 Supplemental Experiments
Our experiments were conducted using three RTX 2080 Ti GPUs, each with 11GB of memory.